For two consecutive integers, the sum of the smaller and twice the larger is 29. Find the smaller integer.
Found 2 solutions by JBarnum, richwmiller:
Answer by JBarnum(2146)

x and x+1
x+2(x+1)=29
x+2x+2=29
3x=27
x=9
x+1=10

Answer by richwmiller(17219)

You were close to the right topic.
n+2*(n+1)=29
Solve for n the smaller integer
3n+2=29
3n=27
n=9
n+1=10
check
9+20=29
ok